### v1

# The MUSE eXtremely Deep Field: Detections of circumgalactic Siāii*Ā emission at ā ā thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

Abstract
Context. The circumgalactic medium (CGM) serves as a baryon reservoir that connects galaxies to the intergalactic medium and fuels star formation. The spatial distribution of the metal-enriched cool CGM has not yet been directly revealed at cosmic noon (ā), as bright emission lines at these redshifts are not covered by optical integral field units.
Aims. To remedy this situation, we aim for the first-ever detections and exploration of extended Siāii*Ā emission (low-ionization state, LIS), referred to as āSiāii*Ā halosā, at redshifts ranging from to as a means to trace the metal-enriched cool CGM.
Methods. We use a sample of 39 galaxies with systemic redshifts of ā measured with the [Cāiii] doublet in the MUSE Hubble Ultra Deep Field catalog, which contains integration times spanning from to hours. We search for extended Siāii*1265, 1309, 1533 emission (fluorescent lines) around individual galaxies. We also stack a subsample of 14 UV-bright galaxies.
Results. We report five individual detections of Siāii* halos. We also confirm the presence of Siāii* halos in stacks for the subsample containing UV-bright sources. The other lines do not show secure detections of extended emission in either individual or stacking analyses. These detections may imply that the presence of metal-enriched CGM is a common characteristic for UV-bright galaxies. To investigate whether the origin of Siāii*Ā is continuum pumping as suggested in previous studies, we check the consistency of the equivalent width (EW) of Siāii*Ā emission and the EW of SiāiiĀ absorption for the individual halo object with the most reliable detection. We confirm the equivalence, suggesting that photon conservation works for this object and pointing toward continuum pumping as the source of Siāii*. We also investigate Siāii*Ā lines in a RAMSES-RT zoom-in simulation including continuum pumping and find ubiquitous presence of extended halos.
Key Words.:
Galaxies: high-redshift ā galaxies: formation ā galaxies: evolution ā galaxies: halos1 Introduction
The circumgalactic medium (CGM) is the baryonic matter that envelops galaxies. The most common definition of the CGM is the gas outside the interstellar medium (ISM) and inside the virial radius of the host dark matter halo. This region serves as a dynamic interface connecting galaxies to the intergalactic medium (IGM). Metals are ejected from galaxies to the CGM via outflows powered by supernovae explosions, stellar winds, and active galactic nuclei (AGNs). A part of those metals may be mixed with inflowing pristine and recycled gas and accrete onto the galaxies. Hence, the CGM is a gas and metal reservoir of a galaxy and plays a pivotal role in galaxy evolution (e.g., Tumlinson et al. 2017; Péroux & Howk 2020).
Historically, the CGM has been observed in absorption imprinted on quasi-stellar object (QSO) spectra (e.g., Wolfe etĀ al. 1986). This method can trace a wide range of column densities and ionization states of various elements. The CGM is found to be a multiphase medium in terms of its density, temperature, ionization, kinematics, metallicity, and structure (e.g., Werk etĀ al. 2016; Steidel etĀ al. 2016; Chen etĀ al. 2020; Schroetter etĀ al. 2021). However, this method is limited to the lines of sight of background sources. Large programs with recent or upcoming instruments targeting bright galaxies as background sources (tomographic mapping) can reach the IGM scale. For instance, The COSMOS Lyman-Alpha Mapping And Tomography Observations (CLAMATO) survey with Keck/LRIS (Lee etĀ al. 2014) and the Prime Focus Spectrograph-Subaru strategic program (PFS SSP) Galaxy Evolution survey (Greene etĀ al. 2022) have a few Mpc transverse resolution at . Higher spatial resolutions can be achieved using QSO pairs (e.g., Tytler etĀ al. 2009; UrbanoĀ Stawinski etĀ al. 2023), gravitationally lensed QSOs (e.g., Chen etĀ al. 2014; Rubin etĀ al. 2018), and gravitational-arcs as background sources (e.g., Lopez etĀ al. 2018). A galaxy-centered stacking approach is also useful to statistically map gas and metals with impact parameters ranging from kpc to a few Mpc (e.g., Rakic etĀ al. 2012; Turner etĀ al. 2014; Dutta etĀ al. 2024). Recently, Bordoloi etĀ al. (2023) achieves the impact parameters of ā kpc individually, using James Web Space Telescope (JWST)/NIRCam slitlesss grism spectroscopy. These studies enable to reach an outer-CGM scale, but an inner-CGM scale is difficult to investigate. Moreover, the identification of the host galaxies that are responsible for the absorption lines in QSO spectra is a challenging task. Hence, this method cannot provide the spatial distribution of gas and metals around individual host galaxies on an inner CGM scale.
A complementary method to absorption-line studies is direct observations of the CGM in emission, which has been challenging because of its low surface brightness (SBs). Modern optical integral field units (IFUs) such as the Multi-Unit Spectroscopic Explorer (MUSE; Bacon etĀ al. 2010) and the Keck Cosmic Web Imager (KCWI; Morrissey 2018) make it possible to detect diffuse CGM emission around individual host galaxies (e.g., Wisotzki etĀ al. 2016). In particular, extended Ly emission, the so-called āLy haloā, which traces cool and warm hydrogen CGM (e.g., Steidel etĀ al. 2011; Momose etĀ al. 2014; Guo etĀ al. 2023c) has been intensively detected and studied around individual Ly emitters (LAEs) at ā (e.g., Leclercq etĀ al. 2017, 2020; Chen etĀ al. 2021; Claeyssens etĀ al. 2022; Erb etĀ al. 2023). Their diverse profiles and statistical trends as well as origins can now be discussed. Recently, a high fraction of Ly halos around UV-continuum selected (not based on Ly) galaxies at - is reported (Kusakabe etĀ al. 2022). The ubiquitous presence of resorvoirs of hydrogen in the CGM is directly confirmed at cosmic noon and earlier epochs ( with MUSE).
However, metals in the CGM at cosmic noon (ā), which is the peak of star-forming activity of galaxies, have not been directly confirmed in emission, in particular for the cool phase. The CGM is known to be metal enriched even at ā from transverse absorption-line studies (e.g., Turner etĀ al. 2014; Lehner etĀ al. 2016; MĆ©ndez-HernĆ”ndez etĀ al. 2022; UrbanoĀ Stawinski etĀ al. 2023; Bordoloi etĀ al. 2023). This is also suggested by outflowing metals identified by down-the-barrel analysis (e.g., Shapley etĀ al. 2003; Steidel etĀ al. 2010; Du etĀ al. 2018; Sugahara etĀ al. 2017) and by broad-line components seen in spectra (e.g., Carniani etĀ al. 2023; Xu etĀ al. 2023). A popular emission tracer of cool metal-enriched CGM, MgāiiĀ 2796, 2803, is useful only at because at higher redshifts the observed wavelength shifts into the near-infrared and the cosmic dimming effect is more severe (e.g., Rubin etĀ al. 2011; Erb etĀ al. 2012; Martin etĀ al. 2013; Burchett etĀ al. 2021; Zabl etĀ al. 2021; Leclercq etĀ al. 2022; Dutta etĀ al. 2023; Guo etĀ al. 2023a, Pessa et al. 2024 in prep.). Other possible tracers are [Oāii] (e.g., Yuma etĀ al. 2013, 2017; Epinat etĀ al. 2018; Johnson etĀ al. 2018) and very faint Feāii*, , , and (e.g., Finley etĀ al. 2017; Shaban etĀ al. 2022). At ā, [Cāii]158m can be covered by sensitive ALMA bands 6 and 7 ([Cāii]Ā halos; e.g., Fujimoto etĀ al. 2019; Ginolfi etĀ al. 2020). At ā, no bright emission tracer of the cool metal-enriched CGM is available in the observed optical regime. While extremely faint, the most promising tracers are the emission lines Siāii*1265, 1309, 1533, which are not contaminated by other line features at nearby wavelengths given the spectral resolution of current facilities (cf. Cāii*1335 emission and Cāii1334 absorption, which cannot be resolved by the resolving power of MUSE, 3000).
Siāii*Ā are fine-structure fluorescent emission lines from singly ionized silicon, which is a low-ionization state (LIS), with an ionization energy close to HāiĀ (ionization energy for Si and Si+ are 8.1Ā eV and 16.4Ā eV, respectively). The energy levels of the Si+ ion for the three transitions this paper focuses on are shown in Figure 1. The origin of Siāii*Ā is suggested to be ācontinuum pumpingā (or ācontinuum fluorescenceā) rather than collisional excitation or recombination (Shapley etĀ al. 2003). Collisional excitations are dominant only in dense environments with high electron densities, which is higher than typical values in Hāii regions (Shapley etĀ al. 2003). And the recombination rates of Si2+ into the excited Si+ is of the same order as the collisional excitation rates when Si+ and Si2+ have comparable abundance with K (Shull & van Steenberg 1982; Shapley etĀ al. 2003). Shapley etĀ al. (2003) model observed nebular emission lines using the photoionization code CLOUDY (Ferland etĀ al. 1998). They find that any model that reproduces the line ratios of the other lines ([Oāiii], [Oāii], H, Oāiii], and Cāiii]) predict more than one order of magnitude weaker Siāii*Ā emission by collisions and recombination. Therefore, those two processes are unlikely to be the dominant origin of Siāii*Ā emission in Hāii regions. As for the continuum pumping scenario, first, Si+ in neutral clouds absorbs the UV continuum and creates SiāiiĀ resonant absorption features. Second, a fraction of Si+ deexcites to the fine-structure level and emits Siāii*Ā photons (see Figure 1). Siāii*Ā lines are escape channels of resonant transitions and are expected to be as spatially extended as the resonant emission lines.
Fluorescent lines are extremely faint, and Siāii*Ā emission is ā times weaker than Ly emission of LAEs (Steidel etĀ al. 2018). In fact, extended Siāii*Ā emission has not been directly detected. Wang etĀ al. (2020) found weak Siāii*Ā emission lines compared to the associated absorption line in HST/Cosmic Origins Spectrograph (COS) data for 5 galaxies at , which implies that the bulk of Siāii*Ā emission arises on larger scales than the COS aperture. Gazagnes etĀ al. (2023) compare COS spectra of Siāii*1265 for 45 local galaxies (COS Legacy Archive Spectroscopic SurveY, CLASSY) with mock spectra from zoom-in simulations (Mauerhofer etĀ al. 2021). They find stronger Siāii*Ā emission lines in the simulations than observed, and they argue that aperture losses of COS can explain the weakness of observed Siāii*Ā and that Siāii*1265 is spatially extended. Very recently, Vasan G.Ā C. etĀ al. (2024) study spatially-resolved outflow properties for a lensed star-forming galaxy at with KCWI and show that Siāii*Ā emission is more extended than the continuum for a certain direction extracted with a pseudo slit.
In this project, we search for Siāii*Ā halos around galaxies at cosmic noon (at ā) as a tracer of metal-enriched cool CGM using integral field unit data to remedy the problem. We use data from the MUSE Hubble Ultra Deep Field (HUDF) survey (Bacon etĀ al. 2023) and test the presence of Siāii*Ā halos with surface brightness profiles of Siāii*Ā and UV continuum. We report the first detections of Siāii*Ā halos. We also stack the MUSE data for a UV-bright subsample to study the general extent of Siāii*Ā and compare it with simulations.
The paper is organized as follows. In Section 2, we describe the data and the sample construction. Section 3 presents methods and results of the search for individual Siāii*Ā halos and those of a stacked subsample. In Section 4, we discuss photon conservation for Siāii, compare observed results with those of zoom-in simulations, and discuss how to increase a sample size with different selection criteria. Finally, conclusions are given in Section 5. Throughout this paper, we assume the Planck 2018 cosmological model (Planck Collaboration etĀ al. 2020) with a matter density of , a dark energy density of , and a Hubble constant of km s-1 Mpc-1 (). Magnitudes are given in the AB system (Oke & Gunn 1983). All distances are in physical units (kpc), unless otherwise stated.
2 Data and sample
To search for extended Siāii*Ā emission of galaxies, we utilize deep MUSE data from data release 2 (DR2) of the MUSE HUDF surveys (Bacon etĀ al. 2023), where rich multi-wavelength data are also available (see Section 2.1). We construct our sample with the DR2 catalog using the systemic redshifts for the individual search and stacking analyses (see Sections 2.2 and 2.3). We create masks to avoid potential contaminations from neighboring galaxies (see Section 2.4).
2.1 Data and catalogs
The MUSE HUDF DR2 data are obtained as a part of the MUSE guaranteed time observations (GTO) program described in Bacon etĀ al. (2023). We use two fields among the MUSE HUDF DR2 data set: a single arcmin2 pointing with a 31-hour depth (udf-10, see also Bacon etĀ al. 2017 and Inami etĀ al. 2017), and the adaptive-optics (AO) assisted MUSE eXtremely Deep Field (mxdf), whose maximum integration time is 141 hours with a 1-arcmin diameter field of view. They are located inside the Hubble eXtreme Deep Field (XDF; Illingworth etĀ al. 2013) with deep HST data (see Bacon etĀ al. 2023, for more details). The udf-10Ā covers the optical wavelength range from 4750Ā Ć Ā to 9350Ā Ć , while mxdfĀ covers from 4700Ā Ć Ā to 9350Ā Ć , with an AO gap from 5800Ā Ć Ā to 5966.25Ā Ć . The spectral resolving power of MUSE varies from to 3750 at 4700Ā Ć Ā to 9350Ā Ć , respectively. The full width at half maximums (FWHMs) of the Moffat point spread function (Moffat PSF, Moffat 1969) are () at the blue wavelength edge and () at the red wavelength edge for the mxdfĀ (udf-10). The point-source flux limit for an unresolved emission line in the mxdfĀ (the udf-10) is erg s-1 cm-2Ā ( erg s-1 cm-2), at around 7000Ā Ć Ā (not affected by OH sky emission), which corresponds to a surface brightness limit for an unresolved emission line of erg s-1 cm-2 arcsec-2Ā ( erg s-1 cm-2 arcsec-2, see Bacon etĀ al. 2023).
The MUSE HUDF DR2 catalog includes 2221 sources from to , which are selected by emission-line detections as well as through HST priors. The details of the catalog construction are described in Bacon etĀ al. (2023). Each MUSE object has a source file, which is in the format of MPDAF111https://mpdaf.readthedocs.io/en/latest/ multi-fits format and is composed of various data files such as minicubes, images, and spectra. It is available on the AMUSED website222https://amused.univ-lyon1.fr/. We use 5ā5ā minicubes from each of the source files. We create continuum-subtracted minicubes following the method used in Kusakabe etĀ al. (2022), taking the median for each pixel within a spectral window of 100 slices (masking km s-1 around the Ly wavelength if the spectrum covers). The continuum subtraction is useful not only to investigate emission lines, but also to remove neighboring sources around a target source. In order to identify HST counterparts, calculate UV magnitudes, and create masks, we use the HST catalog and data from Rafelski etĀ al. (2015).
2.2 Sample selection
We construct a sample of galaxies at ā, for which the [Cāiii]1907, Cāiii]1909 doublet nebular lines and at least one of the Siāii*1265, 1309, 1533 lines are redshifted into the MUSE wavelength range. The available redshift ranges for the three Siāii*Ā lines are shown in the top panel in Figure 2. We require sources to have a secure spectroscopic redshift with and a Cāiii] signal-to-noise ratio (SNR) of . We visually inspect Cāiii] in spectra and images extracted from MUSE cubes to ensure these criteria. We use systemic redshifts measured by the Cāiii] line rather than reference redshifts (REFZs) in the DR2 catalog, which are given by Ly at or absorption lines at . Siāii*Ā lines are expected to be located at the systemic redshifts (e.g., France etĀ al. 2010; Dessauges-Zavadsky etĀ al. 2010; Jaskot & Oey 2014; Wang etĀ al. 2020), though redshifted Siāii*Ā lines are also reported, which could be due to the contamination of the absorption at the blue edge for low-resolution spectra or stacking of spectra for different sources (e.g., Shapley etĀ al. 2003; Erb etĀ al. 2010; Berry etĀ al. 2012). As Siāii*Ā lines are extremely faint, secure measurements of systemic redshifts enable us to not rely on the presence of Siāii*Ā in 1D spectra on a galaxy scale and to directly search for spatially extended Siāii*Ā in narrowband images extracted from MUSE cubes. We limit our sample to galaxies with a secure and isolated HST counterpart (based on Rafelski etĀ al. 2015, MAG_SRCRAF) with a high HST matching confidence level of MCONF in the DR2 catalog. In total, we have 39 galaxies, whose absolute UV magnitudes () and systemic redshifts () are shown in the bottom panel of Figure 2. The is calculated at rest-frame 1600 Ā Ć Ā by fitting 2 or 3 HST bands with a power-low model (with parameters being an amplitude and a UV slope ).
We would like to note that our sample is constructed simply with the deepest data for this pilot study. We discuss possible strategies to extend the sample in Section 4.3.
2.3 Stacked sample
The subsample used for the stacking analysis is restricted to bright UV-continuum galaxies using an apparent magnitude cut of , in order to enhance the SNR of the stacked images (see Section 3.3 and Figure 2). The origin of Siāii*Ā is predicted to be continuum pumping as mentioned in Section 1, and a bright UV continuum is necessary to have a bright Siāii*Ā emission line. Indeed, Siāii*Ā emission lines are stronger in 1D stacked spectra for subsamples of MUSE LAEs with brighter UV continuum (such as those with higher stellar masses and with brighter UV magnitudes) than the subsample counterparts in Feltre etĀ al. (2020). The number of sources for the stacking subsample is 14. We would like to note that applying fainter magnitude cuts indeed reduces the SNRs of the Siāii*Ā lines.
2.4 Masks
In order to exclude pixels that might be affected by bright neighboring objects in Siāii*Ā narrowband images extracted from the minicubes, we create neighboring object masks. First, we define target pixels. We use an HST segmentation map and mask pixels not corresponding to a main target on a 5ā5ā HST/F775W cutout for each object. It is convolved with the MUSE moffat PSF at the redshifted Siāii*Ā wavelength for the main target and rebinned to match the MUSE pixel scale. Then, we apply a threshold value of 0.1 to the peak-normalized convolved cutout to delimit the spaxels of the target. Lower threshold values such as 0.05 cause contamination of neighboring objects inside the target spaxels. Using a higher threshold value of 0.2 does not change the main results in this paper but enhances the contaminated pixels outside the following neighboring object masks. Therefore, we adopt the value 0.1.
Second, we use a 5ā5ā HST/F775W cutout whose sky and the main target are masked with the HST segmentation map for each source. This cutout is convolved with the MUSE PSF and rebinned to match the MUSE pixel scale. Then, we define pixels brighter than a certain threshold on the cutout, except for the target pixels, as masked pixels. The applied threshold value is half of the value above in the absolute sense (without normalization).
3 Results
3.1 Extracted images and surface brightness profiles
For each source, we create a Siāii*Ā emission narrowband (NB) image from the continuum-subtracted minicube. We sum fluxes in a window of 200 km s-1 around the Siāii*Ā wavelength, excluding NaN (masked) spaxels, wavelength slices affected by OH skylines, and the AO gap in the cube. The variance image of the NB is created from that of the cube with error propagation. A UV continuum broadband (BB) is created from the median filtered original minicube with a window of 300 wavelength slices (375Ā Ć ) around the Siāii*Ā wavelength (masking below Ā kmĀ s-1 from the Ly wavelength) to obtain a good SNR. We also exclude NaN in the cube and the AO gap. If the wavelength window is not fully covered, we use as many spectral slices as possible. The variance image of the BB is assumed to be identical to that of the mean-filtered BB created from the same cube with error propagation. We apply a neighboring object mask (see Section 2.4) and measure SB profiles of these two bands using PHOTUTILS (Bradley etĀ al. 2021). The aperture and annuli are centered at the HST counterpart position with 1-pixel () annuli from to and then 2-pixel annuli at and , followed by a 3.5-pixel annulus at , to enhance the SNR at outer radii. The azimuthally-averaged radial SB profiles are calculated from the effective areas and the measured flux above. We normalize the SB profile of the UV continuum at the center () to that of Siāii*.
The top panels in Figure 3 show the Siāii*Ā emission NB, the UV-continuum BB, and the SB profile for a highlighted object, MID. The details of this object are explained later (Section 3.2).
3.2 Identification of Siāii*Ā halos
To identify extended Siāii*Ā emission, we compare the shape of the azimuthally-averaged radial SB profiles of the UV continuum and the Siāii*Ā emission. We take account of uncertainties in SB profiles of Siāii*Ā emission only (blue shades in Figure 3) as those of UV continuum are negligibly small (gray shades). If more than two adjacent data points of the Siāii*Ā SB profiles deviate by more than from the normalized continuum, we identify it as extended Siāii*. The significance of the presence of the halo with this criterion corresponds to more than 97.5%333The probability for two adjacent data points with , which is caused by chance due to noises (i.e., fake detections), is =2.5% with an assumption that noises follow the normal distribution.. We would like to note that, as we use circular photometry, our test may miss non-circular halos as discussed in Kusakabe etĀ al. (2022) for Ly halos. Our halo search is not complete in that sense but secure.
Among the 39 sources, we detect five Siāii*Ā halos for the Siāii* 1533 line (MID, , , , ; see Figure 3). These are the first detections of individual Siāii*Ā halos. The Siāii*Ā lines in the DR2 reference spectra show that there are no significant velocity shifts for the line peaks of Siāii*1533 (Figure 3; the systemic redshifts are given by the [Cāiii]1907, Cāiii]1909 doublet nebular lines, see Section 2.2).
Among the five Siāii*Ā halo sources, MID=, shown in the top panels of Figure 3, is the best-case object. The Siāii*Ā emission profile shows three data points that deviate by more than from the normalized continuum. This object is well-isolated and without a spatial offset of Siāii*Ā from the UV continuum. The images are not significantly affected by masks, which makes the SB profile test secure. The Siāii*Ā emission line can be clearly confirmed on the spectrum shown in Figure 3. MID and MID= are a pair of galaxies, which could contaminate each otherās 5ā 5ā images despite the neighboring object mask. Moreover, MID= is an X-ray detected AGN (Luo etĀ al. 2017), whose Siāii*Ā halo might have a different origin from other sources. The Siāii*Ā NB of MID= shows a peak offset from the UV center measured with the HST data, with two neighboring LAEs in the DR2 catalog, which spatially overlap with the extended Siāii*Ā emission. The Siāii*Ā SB of MID= is noisy even though it meets our halo criterion. Considering these details on the five sources, we highlight MID= in the following sections.
The redshifts of the individually-detected Siāii*1533 halos range from 2.23 to 3.32 with four of them at . The non detections at could be due to the high noise levels of MUSE data at red wavelengths (severe night sky and low sensitivity; Bacon etĀ al. 2023) and the cosmic dimming effect on the SB profiles (by a factor of 2.7 between and ).
Another interesting result is that we cannot confirm extended emission for any of the other Siāii*Ā lines (1265, 1309). In the optically thin limit, the strengths of fluorescent lines are expected to depend on the strength of the paired absorption lines, which are determined by their oscillator strengths, . In particular, the Siāii resonant absorption line corresponding to the Siāii*1265 line has a ten times larger than that of the Siāii resonant absorption line corresponding to the Siāii*1533. However, in the simulations of Mauerhofer etĀ al. (2021) there is no significant difference in the equivalent widths of Siāii*1265 and Siāii*1533 absorption lines because both lines are saturated (Mauerhofer et al. in prep.). In such an optically thick regime, the equivalent widths for the corresponding fluorescent emission (Siāii*1265 and Siāii*1533) are similar in the simulations. Therefore, non detections of Siāii*1265 as well as Siāii*1309 would not be due to fainter total fluxes for these lines than for the 1533 line. The reason for nondetections of Siāii*1265 and Siāii*1309 halos are not clear. The samples for Siāii*1265, 1309 lines are limited to high redshifts ( and , respectively), and none of them are bright in apparent UV magnitudes, which are directly related to SNRs (see Figure 2). Bright sources would be crucial for the detection of diffuse Siāii*Ā lines if the origin of Siāii*Ā is continuum pumping. However, we cannot distinguish whether it is caused by differences in lines, samples, or noise levels. We note that the Siāii* 1533 lines are also brighter than the other two lines in the stacked MUSE LAE spectra of Feltre etĀ al. (2020) and Kramarenko etĀ al. (2023), which might also be due to the difference in objects stacked for different Siāii*Ā lines.
One of the individually-detected objects, MID= at , has 3 Siāii*Ā lines covered with MUSE. We coadd the NBs and BBs for the 3 Siāii*Ā lines by taking the median (with the same method as used in Section 3.3) and show them in the fifth row of Figure 3. Extended Siāii*Ā is confirmed with a better SNR, which implies the Siāii*1265, 1309 lines are also spatially extended.
3.3 Stacking analysis
To test the general presence of Siāii*Ā halos, we stack 14 UV-bright sources (see Section 2.3). The numbers of sources with the magnitude cut of available for each of the Siāii*1265, 1309, 1533 lines are five, four, and 13, respectively, which are limited by the MUSE wavelength coverage and the AO gap (see Figure 2)444The MIDs for the Siāii*1265 subsample are 50, 76, 6666, 6700, and 8048. Those for the Siāii*1309 are 50, 76, 6700, and 8048. Those for the Siāii*1533 are 24, 35, 50, 51, 76, 1141, 1316, 1395, 6664, 6666, 6700, 7600, and 8513.. Therefore, we only stack a subsample for the Siāii*1533 line and take the mean and median of the Siāii*Ā NBs and BBs without weighting, while masking neighboring objects. We also stack 14 sources by combining the NBs and BBs for the three lines. The variance images for the mean stacking, which we use for our halo test, are created based on error propagation. Then, we apply the SB profile test for the mean-stacked images with the same method as used in Sections 3.1 and 3.2, as the uncertainties can be calculated correctly for the mean but not for the median.
The results of this stacking experiment are shown in Figure 4. As shown in the top row, we confirm the existence of a Siāii*1533 halo with five data points that deviate by more than from the normalized continuum profile. The median profile of Siāii*1533 is consistent with the mean profile and deviates from the median continuum as well, in particular at large radii. These tests may suggest that the detected halo in the mean profile is not dominated by a few extreme objects. This is the first secure detection of a stacked Siāii*Ā halo. It may imply that the individually detected sources are not extreme cases, Siāii*Ā halos must be common for UV-bright galaxies at ā.
The bottom row of Figure 4 shows the results of stacking all the Siāii*Ā narrowbands and corresponding UV continuum bands for the 14 sources (three lines; in total 22 images). We find that the mean-stacked Siāii*Ā deviates by more than from the mean-stacked UV continuum for six data points. The median-stacked Siāii*Ā is consistent with the mean-stacked Siāii*, and the deviation from the normalized continuum is similarly seen with the median-stacked profiles.
We note that removing MID, which is an X-ray detected AGN with extended Siāii*Ā (see Section 3.2), does not change the results and that the other 13 sources do not have a counterpart in the deep X-ray catalog (Luo etĀ al. 2017). We also note that stacking without individually-detected Siāii*Ā halos results in nondetections of extended Siāii*. As shown in Figure 2, all the individually-detected sources have a bright (and also a bright ) and are located at relatively low redshifts. It is not surprising that removing such five sources results in nondetections. If the origin of the Siāii*Ā is continuum pumping, as we expect, it then should be more challenging to detect Siāii*Ā halos without these bright sources. We also stack various subsamples including all the sources without individually-detected halos and get nondetections of extended Siāii*Ā emission. Future instruments such as BlueMUSE (Richard etĀ al. 2019) or a larger sample compiling MUSE and KCWI archive data will allow us to investigate the dependence of Siāii*Ā halos on sample properties such as (see Section 4.3). It will enable us to draw a firm conclusion of whether Siāii*Ā halos are indeed common for UV-bright galaxies at ā or not.
4 Discussion
4.1 Testing photon conservation
Shapley etĀ al. (2003) suggest that the origin of the Siāii*Ā is continuum pumping (see Section 1 for other origins). For the Siāii* line, Si+ in neutral clouds makes a Siāii resonant absorption line, and then a fraction of Si+ deexcites to the fine-structure level, emitting SiāiiĀ fluorescence photons at 1533Ā Ć Ā (see Figure 1). As the SiāiiĀ is a resonant line, the Siāii*Ā line can be regarded as an escape channel of photons from resonant scattering. If dust extinction is negligible and the escape is isotropic, then the equivalent width (EW) of Siāii*Ā emission and that of the corresponding resonant absorption should be comparable because of photon conservation. However, it is possible for the absorption and emission to have different EWs for one given direction of observations even if the origin is continuum pumping without dust, because of the complex gas distribution in front of stars in galaxies (e.g., Prochaska etĀ al. 2011b; Carr etĀ al. 2018). Simulations predict that absorption lines depend sensitively on the direction of observations (Mauerhofer etĀ al. 2021; Gazagnes etĀ al. 2023). Below, we assume an isotropic case.
To investigate the scenario of continuum pumping, we use the isolated, high-SN object MID (see Section 3.2) and test photon conservation between SiāiiĀ absorption and associated Siāii*Ā emission, which is commonly assessed with the EWs of these lines (e.g., Prochaska etĀ al. 2011a; Wang etĀ al. 2020). First, we check the flux curve of growth (CoG) for the continuum and the emission line to determine the spatial extent. Second, we measure the EW curve of growth for Siāii*Ā emission and compare it with the absorption EW. We measure EWs from the MUSE data (without HST), which makes the EW comparison fair. Ideally, we should measure the curve of growth for the absorption. However, due to the limited depth and spatial resolution of our data, we have to assume that the absorption follows the same spatial profile as that of the continuum. The left panel of Figure 5 shows normalized fluxes within growing apertures for continuum (black) and Siāii* (blue). Fluxes are measured from the Siāii* NB image and the corresponding UV continuum BB image (Section 3.1) with PHOTUTILS and the neighboring object mask with corrections of masked areas. On a galaxy-scale of R, the aperture includes about 80% of the continuum but only 50% of the Siāii*. An R aperture captures most of the Siāii*Ā flux.
The right panel shows rest-frame EWs for Siāii* emission within growing apertures (blue), compared with the EW for Siāii absorption measured at R, EWabs(Siāii)Ā Ć Ā (red). Rest-frame EWs are measured in spectra extracted from the original minicube (to which the neighboring object mask is applied) with growing apertures. The continuum flux density is measured by the median filtering each spectrum from 50-slice shorter wavelength (62.5Ć ) than the absorption to 50-slice longer wavelength than the emission. Absorption and emission fluxes are measured as differences from the continuum. The EW for Siāii* is EWem(Siāii*)Ā Ć Ā at R, which is smaller than EWabs(Siāii)Ā Ć . EWem(Siāii*) increases with R and reaches EWem(Siāii*)Ā Ć Ā at R, which is consistent with EWabs(Siāii) at R within the 1 uncertainties (see Figure 6 for the spectra), although the EWs have large measurement errors. Therefore, the absorption EW on the galaxy scale and the fluorescent emission EW on the CGM scale are consistent with each other. This is compatible with photon conservation on the CGM scale, albeit with large uncertainties. It implies that the origin of extended Siāii*Ā emission can be the continuum pumping as predicted under the assumption of isotropic escape and no dust. In that case, either a small amount of dust, a low optical depth, or a combination of both are implied. Unfortunately, data for the other four sources with individual Siāii*Ā halos are either noisy or have accompanying galaxies inside their 5ā5ā minicubes, which prevents us from obtaining a firm conclusion about the mechanisms for the four sources. The trend of increasing EW of Siāii*Ā emission with radius beyond the galaxy scale is predicted by simulations (Mauerhofer etĀ al. 2021), which are compared with slit spectroscopy of low-redshift galaxies in Gazagnes etĀ al. (2023, see also ).
4.2 Comparisons with zoom simulations
We compare our stacked Siāii* profile with those from the cosmological zoom-in simulations of Mauerhofer etĀ al. (2021), which are used in Gazagnes etĀ al. (2023) and Blaizot etĀ al. (2023). Mauerhofer etĀ al. (2021) present three snapshots containing a relatively small galaxy (such as MUSE LAEs) at , , and . The stellar mass for the galaxy is Ā Mā, and the star formation rate ranges from to 5Ā MāĀ yr-1 for the three snapshots. Siāii*photons in this simulations originate from continuum pumping. We provide mock MUSE cubes for observations from 12 directions for each snapshot in the rest frame. This original mock cube has a size of pixels corresponding to Ā Ć Ā (from 1521.83Ā Ć Ā to 1538.42Ā Ć ) with a spaxel scale of /pix. Then, to match the size of observed and simulated galaxies, we measure the median size of the simulated galaxy among the 36 directions. A UV-continuum BB image is created from each median-filtered original cube with a window from 1528.5Ā Ć Ā to 1531.5Ā Ć Ā (the continuum between SiāiiĀ absorption and Siāii*Ā emission). We define the galaxy center by identifying the brightest pixel, as in Gazagnes etĀ al. (2023), for each UV continuum BB image after HST PSF convolution (F606W in Rafelski etĀ al. 2015). The half-light radius (R50) is measured around the galaxy center with the original mock cube (without HST PSF convolution) using PHOTUTILS. The median R50 for the simulated galaxy is (3.5 pix). This is 1.7 times smaller than the median R50 for the 14 stacked galaxies (R50= with F606W after PSF correction). Since we do not have a simulated galaxy that is large enough in Mauerhofer etĀ al. (2021), we manually rescale the original mock cube by a factor of 1.7 by interpreting the spaxel scale as /pix, which is referred to as an intrinsic mock cube below. The intrinsic mock cubes are cut out around the galaxy center defined above and are 5 arcsec each. The 5ā cubes are convolved with the MUSE PSF and line spread function (LSF) described in Bacon etĀ al. (2023) at the median redshift of the stacked sample (). Then the convolved mock cubes are rebinned to match the MUSE cubes in the rest frame. We apply similar analyses to the 36 mock MUSE cubes as used for the observational data (see Section 3). A slight difference from the method for the MUSE data is continuum subtraction. Continuum cubes used for continuum subtraction are created using median filtering from 1528.5Ā Ć Ā to 1531.5Ā Ć ā which are also used for the UV-continuum BB images. The 1 uncertainties for stacked mock data are derived from 15.87 to 84.12 percentiles of SB profiles among individual mocks. We would like to note that the simulations assume cosmological parameters of km s-1 Mpc-1 and , which should not have a significant impact on the results below. We convert distances in the unit of arcsec for the simulated data to those in the unit of kpc at with the Planck 2018 cosmological model (Planck Collaboration etĀ al. 2020) as the galaxy sizes are matched and MUSE PSF are convolved in the unit of arcsec.
The left panel of Figure 7 shows a comparison of the observed and simulated mean-stacked SB profiles. The simulated UV continuum (gray dotted line) is well matched with the observations (black solid line), thanks to our manual 1.7 times expansion. The simulated Siāii* profile (green dotted line) is more extended than the simulated UV continuum (gray dotted line) and consistent with the observed Siāii* (blue solid line), though simulated Siāii*Ā halo is slightly less extended than observed within uncertainties.
The mean profiles can be biased toward bright halo sources, so we check whether the mean-stacked profiles are consistent with the median-stacked profiles for the Siāii*Ā and the continuum (green and gray crosses, respectively) in the right panel of Figure 7. We confirm that they are consistent with each other, as for the MUSE observations of Siāii*. The right panel also compares simulated mean-stacked SB profiles with PSF and LSF convolution versus without convolution, indicated by dotted and solid lines, respectively. The simulated galaxy indeed has significantly more extended Siāii* than the UV continuum in the intrinsic mock cubes, but the difference is mostly hidden by the MUSE PSF convolution555We check individual mocks and find that the shapes of the individual SB profiles (green dotted line) are fully dominated by the PSF..
We also check EW CoG of Siāii* emission and Siāii absorption lines for the simulations as shown in the left panel of Figure 8. The EWs of Siāii* emission and Siāii absorption are consistent at a CGM scale for both intrinsic cubes and PSF-convolved cubes, which are also seen on the mock spectra (see the middle and right panels of Figure 8). It means that the photon conservation works for the simulations with continuum pumping in the sense that the scattered photons are not more attenuated by dust than that stellar continuum. This is consistent with our observations for MID=1141.
From these comparisons, we conclude that our simulations with continuum pumping scenario can reproduce the observations. This lends extra support for the interpretation of Siāii*Ā emission as a signature of continuum pumping. We note that the selection bias with the UV magnitude cut may have an impact on the observed result and this comparison. The mock observations are re-scaled by a factor of to correct for the fact that the simulated galaxy is smaller than the observed ones (see above). The simulated galaxy is fainter than the observed one, though the range of absolute magnitudes overlaps ( from -18 to -19). It would be interesting to test the selection bias with a larger and brighter simulated galaxy, or several, maybe at a slightly lower redshift.
4.3 Possible extension of the sample
We discuss possible strategies to extend the sample with existing data in this section. There are a few ways to increase the sample size: (1) sources at , which have other nebular line detections such as the Oāiii]1661, 1666 doublet and Heāii in the same catalog and the same field (with the same integration time), (2) sources located in shallower fields such as mosaicĀ field (99-pointing with 10-hour integration, in the same catalog of Bacon etĀ al. 2023) and MUSE-WIDE field (1-hour integration Herenz etĀ al. 2017; Urrutia etĀ al. 2019; Kerutt etĀ al. 2022), (3) UV-bright sources at the same redshifts in other fields with similarly long integration times ( hours) in MUSE archival data (e.g., Fossati etĀ al. 2019; Lofthouse etĀ al. 2020). The samples (1) and (2) are tested with the MUSE DR2 catalog (not including MUSE WIDE) in the same manner. We find that these less restrictive selection criteria do not help to increase the number of detections of Siāii*Ā halos. Therefore, (3) using deep MUSE archival data for UV-bright galaxies with the same selection criteria would be the best way to extend the sample in the future.
In addition, (4) gravitationally lensed sources observed with MUSE (e.g., Richard etĀ al. 2021; Claeyssens etĀ al. 2022, Claeyssens et al. 2024 in prep.), would be useful to investigate more compact Siāii*Ā halos in the future (see the right panel of Figure 7), which would be hidden by the PSF without the power of the magnification. We would like to note that they are not included in this pilot study as magnification can add complications and uncertainties. To investigate Siāii*Ā halos at higher redshifts (ā), whose Siāii*Ā lines are covered with MUSE, we could use (5) sources with systemic redshifts measured by nebular lines with James Webb Space Telescope (JWST), and (6) MUSE LAEs, whose systemic redshifts can be estimated by empirical relationships based on the Ly line (Verhamme etĀ al. 2018). We do not include these sources in this pilot study, considering uncertainties in wavelength calibration (e.g., Maseda etĀ al. 2023; DāEugenio etĀ al. 2024; Meyer etĀ al. 2024) and more severe cosmic dimming effects. It is future work for us to investigate higher- sources.
5 Conclusions
To study the spatial distribution of the metal-enriched cool CGM at cosmic noon in emission, we search for extended Siāii*Ā emission (fluorescent lines) using the MUSE HUDF data and catalog with 30-140 hour integration times. We construct a sample of 39 galaxies with systemic redshifts at ā, at which the [Cāiii]1907, Cāiii]1909 doublet nebular lines and at least one of the Siāii*1265, 1309, 1533 lines are redshifted into the MUSE wavelength range. Our major results are summarized as follows.
1.
Five individual Siāii* halos are confirmed to be present from statistical tests with surface brightness profiles. These are the first detections of extended Siāii*Ā emission.
2.
We stack images of Siāii*1533 line for a subsample of 13 UV-bright galaxies. We confirm the presence of stacked Siāii* halos, which may imply that metal-enriched CGM is common for UV-bright galaxies. We also stack images of Siāii*1265, 1309, 1533 for 14 UV-bright galaxies and detect a Siāii*Ā halo. If we stack fainter galaxies or remove the five individually-detected halos from stacking, we get nondetections of Siāii*Ā halos. We need a larger sample to draw a firm conclusion.
3.
We find that the EW of the absorption line and fluorescent emission line are roughly equal when integrated out to large distances for MID=1141. This suggests that photons are conserved and that the emission line is mostly due to pumping from the stellar continuum.
4.
We test the presence of Siāii*Ā halos in zoom-in simulations, which account only for continuum pumping. After re-scaling the mock observations to correct for the fact that the simulated galaxy is smaller than the observed one, we find that the simulated halos are consistent with the observed stacked halo within uncertainties. This lends extra support for the interpretation of Siāii*Ā emission as a signature of continuum pumping.
5.
The best way to extend our sample with existing data in the future is using deep MUSE and KCWI archival data for UV-bright galaxies with the same selection criteria.
KCWI or BlueMUSE (Richard etĀ al. 2019) will allow us to investigate HI Ly halos for Siāii*Ā halos detected with MUSE. BlueMUSE will also make it possible to individually detect Siāii*Ā halos of more diverse sources and to study their statistical properties thanks to a more moderate cosmic dimming effect at (by a factor of three, compared to ). Such larger samples would be useful to improve the stacking experiments (see also Section 4.3) and to study spectral variations at different radii from the centers of galaxies as done for Ly (Wisotzki etĀ al. 2018; Guo etĀ al. 2023c, b). Such studies will be advanced more with better sensitivities and higher spatial resolutios achieved by extremely large telescopes such as Thirty-meter telescope (TMT), one of whose key science cases is CGM mapping with WFOS (Skidmore etĀ al. 2015), as well as European-Extremely large telescope (E-ELT).
Acknowledgements.
HK acknowledges support from the Japan Society for the Promotion of Science (JSPS) Overseas Research Fellowship (202160056) as well as JSPS Research Fellowships for Young Scientists (202300224, 23KJ2148). HK thanks Yuri Nagai, an academic assistant at NAOJ, for her wonderful support. VM acknowledges support from the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO) grant 016.VIDI.189.162 (āODINā). AV and TG are supported by the SNF grant PP00P2 211023. T.N. acknowledges support from Australian Research Council Laureate Fellowship FL180100060. I.P. acknowledges funding by the European Research Council through ERC-AdG SPECMAP-CGM, GA 101020943. This work is based on observations taken by VLT, which is operated by European Southern Observatory. This research made use of Astropy666http://www.astropy.org, which is a community-developed core Python package for Astronomy (The Astropy Collaboration etĀ al. 2013, 2018), and other software and packages:MPDAF (Piqueras etĀ al. 2019), PHOTUTILS, Numpy (Harris etĀ al. 2020), Scipy (Virtanen etĀ al. 2020), and matplotlib (Hunter 2007).References
- Bacon etĀ al. (2010) Bacon, R., Accardo, M., Adjali, L., etĀ al. 2010, Proceedings of the SPIE, 7735, 773508, conference Name: Ground-based and Airborne Instrumentation for Astronomy III Place: eprint: arXiv:2211.16795 ADS Bibcode: 2010SPIE.7735E..08B
- Bacon etĀ al. (2023) Bacon, R., Brinchmann, J., Conseil, S., etĀ al. 2023, Astronomy & Astrophysics, 670, A4
- Bacon etĀ al. (2017) Bacon, R., Conseil, S., Mary, D., etĀ al. 2017, Astronomy & Astrophysics, 608, A1
- Berry etĀ al. (2012) Berry, M., Gawiser, E., Guaita, L., etĀ al. 2012, The Astrophysical Journal, 749, 4
- Blaizot etĀ al. (2023) Blaizot, J., Garel, T., Verhamme, A., etĀ al. 2023, Monthly Notices of the Royal Astronomical Society, 523, 3749, aDS Bibcode: 2023MNRAS.523.3749B
- Bordoloi etĀ al. (2023) Bordoloi, R., Simcoe, R.Ā A., Matthee, J., etĀ al. 2023, EIGER IV: The cool 10$^4$K circumgalactic environment of high-$z$ galaxies reveals remarkably efficient IGM enrichment, arXiv:2307.01273 [astro-ph]
- Bradley etĀ al. (2021) Bradley, L., SipÅcz, B., Robitaille, T., etĀ al. 2021, astropy/photutils: 1.3.0
- Burchett etĀ al. (2021) Burchett, J.Ā N., Rubin, K. H.Ā R., Prochaska, J.Ā X., etĀ al. 2021, The Astrophysical Journal, 909, 151
- Carniani etĀ al. (2023) Carniani, S., Venturi, G., Parlanti, E., etĀ al. 2023, JADES: The incidence rate and properties of galactic outflows in low-mass galaxies across 3 < z < 9, publication Title: arXiv e-prints ADS Bibcode: 2023arXiv230611801C
- Carr etĀ al. (2018) Carr, C., Scarlata, C., Panagia, N., & Henry, A. 2018, The Astrophysical Journal, 860, 143
- Chen etĀ al. (2014) Chen, H.-W., Gauthier, J.-R., Sharon, K., etĀ al. 2014, Monthly Notices of the Royal Astronomical Society, 438, 1435
- Chen etĀ al. (2020) Chen, H.-W., Zahedy, F.Ā S., Boettcher, E., etĀ al. 2020, Monthly Notices of the Royal Astronomical Society, 497, 498, publisher: OUP ADS Bibcode: 2020MNRAS.497..498C
- Chen etĀ al. (2021) Chen, Y., Steidel, C.Ā C., Erb, D.Ā K., etĀ al. 2021, Monthly Notices of the Royal Astronomical Society, 508, 19
- Claeyssens etĀ al. (2022) Claeyssens, A., Richard, J., Blaizot, J., etĀ al. 2022, Astronomy & Astrophysics, 666, A78
- Dessauges-Zavadsky etĀ al. (2010) Dessauges-Zavadsky, M., DāOdorico, S., Schaerer, D., etĀ al. 2010, Astronomy and Astrophysics, 510, A26, aDS Bibcode: 2010A&Aā¦510A..26D
- DāEugenio etĀ al. (2024) DāEugenio, F., Cameron, A.Ā J., Scholtz, J., etĀ al. 2024, JADES Data Release 3 ā NIRSpec/MSA spectroscopy for 4,000 galaxies in the GOODS fields, publication Title: arXiv e-prints ADS Bibcode: 2024arXiv240406531D
- Du etĀ al. (2018) Du, X., Shapley, A.Ā E., Reddy, N.Ā A., etĀ al. 2018, The Astrophysical Journal, 860, 75
- Dutta etĀ al. (2023) Dutta, R., Fossati, M., Fumagalli, M., etĀ al. 2023, Monthly Notices of the Royal Astronomical Society, 522, 535
- Dutta etĀ al. (2024) Dutta, S., Muzahid, S., Schaye, J., etĀ al. 2024, Monthly Notices of the Royal Astronomical Society, 528, 3745, aDS Bibcode: 2024MNRAS.528.3745D
- Epinat etĀ al. (2018) Epinat, B., Contini, T., Finley, H., etĀ al. 2018, Astronomy & Astrophysics, 609, A40
- Erb etĀ al. (2023) Erb, D.Ā K., Li, Z., Steidel, C.Ā C., etĀ al. 2023, The Astrophysical Journal, 953, 118, aDS Bibcode: 2023ApJā¦953..118E
- Erb etĀ al. (2010) Erb, D.Ā K., Pettini, M., Shapley, A.Ā E., etĀ al. 2010, The Astrophysical Journal, 719, 1168
- Erb etĀ al. (2012) Erb, D.Ā K., Quider, A.Ā M., Henry, A.Ā L., & Martin, C.Ā L. 2012, The Astrophysical Journal, 759, 26
- Feltre etĀ al. (2020) Feltre, A., Maseda, M.Ā V., Bacon, R., etĀ al. 2020, Astronomy & Astrophysics, 641, A118
- Ferland etĀ al. (1998) Ferland, G.Ā J., Korista, K.Ā T., Verner, D.Ā A., etĀ al. 1998, Publications of the Astronomical Society of the Pacific, 110, 761, publisher: The University of Chicago Press
- Finley et al. (2017) Finley, H., Bouché, N., Contini, T., et al. 2017, Astronomy & Astrophysics, 605, A118
- Fossati etĀ al. (2019) Fossati, M., Fumagalli, M., Lofthouse, E.Ā K., etĀ al. 2019, Monthly Notices of the Royal Astronomical Society, 490, 1451, publisher: OUP ADS Bibcode: 2019MNRAS.490.1451F
- France etĀ al. (2010) France, K., Nell, N., Green, J.Ā C., & Leitherer, C. 2010, The Astrophysical Journal, 722, L80
- Fujimoto etĀ al. (2019) Fujimoto, S., Ouchi, M., Ferrara, A., etĀ al. 2019, The Astrophysical Journal, 887, 107
- Gazagnes etĀ al. (2023) Gazagnes, S., Mauerhofer, V., Berg, D.Ā A., etĀ al. 2023, The Astrophysical Journal, 952, 164, aDS Bibcode: 2023ApJā¦952..164G
- Ginolfi et al. (2020) Ginolfi, M., Jones, G. C., Béthermin, M., et al. 2020, Astronomy & Astrophysics, 633, A90
- Greene etĀ al. (2022) Greene, J., Bezanson, R., Ouchi, M., Silverman, J., & Group, t. P. G. E.Ā W. 2022, The Prime Focus Spectrograph Galaxy Evolution Survey, arXiv:2206.14908 [astro-ph]
- Guo et al. (2023a) Guo, Y., Bacon, R., Bouché, N. F., et al. 2023a, Nature, 624, 53, arXiv:2312.05167 [astro-ph]
- Guo etĀ al. (2023b) Guo, Y., Bacon, R., Wisotzki, L., etĀ al. 2023b, Spatially-resolved Spectroscopic Analysis of Ly$\alpha$ Haloes: Radial Evolution of the Ly$\alpha$ Line Profile out to 60 kpc, arXiv:2309.06311 [astro-ph]
- Guo etĀ al. (2023c) Guo, Y., Bacon, R., Wisotzki, L., etĀ al. 2023c, Median Surface Brightness Profiles of Lyman-$\alpha$ Haloes in the MUSE Extremely Deep Field, arXiv:2309.05513 [astro-ph]
- Harris etĀ al. (2020) Harris, C.Ā R., Millman, K.Ā J., Van DerĀ Walt, S.Ā J., etĀ al. 2020, Nature, 585, 357
- Herenz etĀ al. (2017) Herenz, E.Ā C., Urrutia, T., Wisotzki, L., etĀ al. 2017, Astronomy & Astrophysics, 606, A12
- Hunter (2007) Hunter, J.Ā D. 2007, Computing in Science & Engineering, 9, 90
- Illingworth etĀ al. (2013) Illingworth, G.Ā D., Magee, D., Oesch, P.Ā A., etĀ al. 2013, The Astrophysical Journal Supplement Series, 209, 6
- Inami etĀ al. (2017) Inami, H., Bacon, R., Brinchmann, J., etĀ al. 2017, Astronomy & Astrophysics, 608, A2
- Jaskot & Oey (2014) Jaskot, A.Ā E. & Oey, M.Ā S. 2014, The Astrophysical Journal, 791, L19
- Johnson etĀ al. (2018) Johnson, S.Ā D., Chen, H.-W., Straka, L.Ā A., etĀ al. 2018, The Astrophysical Journal, 869, L1
- Kerutt etĀ al. (2022) Kerutt, J., Wisotzki, L., Verhamme, A., etĀ al. 2022, Astronomy & Astrophysics, 659, A183
- Kramarenko etĀ al. (2023) Kramarenko, I.Ā G., Kerutt, J., Verhamme, A., etĀ al. 2023, Monthly Notices of the Royal Astronomical Society, 527, 9853, arXiv:2305.07044 [astro-ph]
- Kusakabe etĀ al. (2022) Kusakabe, H., Verhamme, A., Blaizot, J., etĀ al. 2022, Astronomy & Astrophysics, 660, A44
- Leclercq etĀ al. (2020) Leclercq, F., Bacon, R., Verhamme, A., etĀ al. 2020, Astronomy & Astrophysics, 635, A82
- Leclercq etĀ al. (2017) Leclercq, F., Bacon, R., Wisotzki, L., etĀ al. 2017, Astronomy & Astrophysics, 608, A8
- Leclercq etĀ al. (2022) Leclercq, F., Verhamme, A., Epinat, B., etĀ al. 2022, Astronomy & Astrophysics, 663, A11, arXiv:2203.05614 [astro-ph]
- Lee etĀ al. (2014) Lee, K.-G., Hennawi, J.Ā F., Stark, C., etĀ al. 2014, The Astrophysical Journal, 795, L12
- Lehner etĀ al. (2016) Lehner, N., OāMeara, J.Ā M., Howk, J.Ā C., Prochaska, J.Ā X., & Fumagalli, M. 2016, The Astrophysical Journal, 833, 283
- Lofthouse etĀ al. (2020) Lofthouse, E.Ā K., Fumagalli, M., Fossati, M., etĀ al. 2020, Monthly Notices of the Royal Astronomical Society, 491, 2057, publisher: OUP ADS Bibcode: 2020MNRAS.491.2057L
- Lopez etĀ al. (2018) Lopez, S., Tejos, N., Ledoux, C., etĀ al. 2018, Nature, 554, 493, aDS Bibcode: 2018Natur.554..493L
- Luo etĀ al. (2017) Luo, B., Brandt, W.Ā N., Xue, Y.Ā Q., etĀ al. 2017, The Astrophysical Journal Supplement Series, 228, 2, publisher: IOP ADS Bibcode: 2017ApJS..228ā¦.2L
- Martin etĀ al. (2013) Martin, C.Ā L., Shapley, A.Ā E., Coil, A.Ā L., etĀ al. 2013, The Astrophysical Journal, 770, 41
- Maseda etĀ al. (2023) Maseda, M.Ā V., Lewis, Z., Matthee, J., etĀ al. 2023, JWST/NIRSpec Measurements of Extremely Low Metallicities in High Equivalent Width Lyman-$\alpha$ Emitters, arXiv:2304.08511 [astro-ph]
- Mauerhofer etĀ al. (2021) Mauerhofer, V., Verhamme, A., Blaizot, J., etĀ al. 2021, Astronomy & Astrophysics, 646, A80
- Meyer etĀ al. (2024) Meyer, R.Ā A., Oesch, P.Ā A., Giovinazzo, E., etĀ al. 2024, JWST FRESCO: a comprehensive census of H$\beta$+[OIII] emitters at 6.8, publication Title: arXiv e-prints ADS Bibcode: 2024arXiv240505111M
- Moffat (1969) Moffat, A. F.Ā J. 1969, Astronomy and Astrophysics, 3, 455, aDS Bibcode: 1969A&Aā¦..3..455M
- Momose etĀ al. (2014) Momose, R., Ouchi, M., Nakajima, K., etĀ al. 2014, Monthly Notices of the Royal Astronomical Society, 442, 110
- Morrissey (2018) Morrissey, P. 2018, The Astrophysical Journal
- Méndez-HernÔndez et al. (2022) Méndez-HernÔndez, H., Cassata, P., Ibar, E., et al. 2022, Astronomy & Astrophysics, 666, A56, arXiv:2206.08923 [astro-ph]
- Oke & Gunn (1983) Oke, J.Ā B. & Gunn, J.Ā E. 1983, The Astrophysical Journal, 266, 713, aDS Bibcode: 1983ApJā¦266..713O
- Piqueras etĀ al. (2019) Piqueras, L., Conseil, S., Shepherd, M., etĀ al. 2019, Astronomical Data Analysis Software and Systems XXVI ASP Conference Series, 521, 545, conference Name: Astronomical Data Analysis Software and Systems XXVI ADS Bibcode: 2019ASPC..521..545P
- Planck Collaboration etĀ al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., etĀ al. 2020, Astronomy & Astrophysics, 641, A6
- Prochaska etĀ al. (2011a) Prochaska, J.Ā X., Kasen, D., & Rubin, K. 2011a, The Astrophysical Journal, 734, 24
- Prochaska etĀ al. (2011b) Prochaska, J.Ā X., Weiner, B., Chen, H.-W., Mulchaey, J., & Cooksey, K. 2011b, The Astrophysical Journal, 740, 91
- Péroux & Howk (2020) Péroux, C. & Howk, J. C. 2020, Annual Review of Astronomy and Astrophysics, 58, 363
- Rafelski etĀ al. (2015) Rafelski, M., Teplitz, H.Ā I., Gardner, J.Ā P., etĀ al. 2015, The Astronomical Journal, 150, 31
- Rakic etĀ al. (2012) Rakic, O., Schaye, J., Steidel, C.Ā C., & Rudie, G.Ā C. 2012, The Astrophysical Journal
- Richard etĀ al. (2019) Richard, J., Bacon, R., Blaizot, J., etĀ al. 2019, BlueMUSE: Project Overview and Science Cases, arXiv:1906.01657 [astro-ph]
- Richard etĀ al. (2021) Richard, J., Claeyssens, A., Lagattuta, D., etĀ al. 2021, Astronomy & Astrophysics, 646, A83
- Rubin etĀ al. (2018) Rubin, K. H.Ā R., OāMeara, J.Ā M., Cooksey, K.Ā L., etĀ al. 2018, The Astrophysical Journal, 859, 146
- Rubin et al. (2011) Rubin, K. H. R., Prochaska, J. X., Ménard, B., et al. 2011, The Astrophysical Journal, 728, 55
- Schroetter et al. (2021) Schroetter, I., Bouché, N. F., Zabl, J., et al. 2021, Monthly Notices of the Royal Astronomical Society, 506, 1355
- Shaban etĀ al. (2022) Shaban, A., Bordoloi, R., Chisholm, J., etĀ al. 2022, The Astrophysical Journal, 936, 77, aDS Bibcode: 2022ApJā¦936ā¦77S
- Shapley etĀ al. (2003) Shapley, A.Ā E., Steidel, C.Ā C., & Pettini, M. 2003, The Astrophysical Journal, 588
- Shull & van Steenberg (1982) Shull, J.Ā M. & van Steenberg, M. 1982, The Astrophysical Journal Supplement Series, 48, 95, publisher: IOP ADS Bibcode: 1982ApJSā¦48ā¦95S
- Skidmore etĀ al. (2015) Skidmore, W., DellāAntonio, I., Fukugawa, M., etĀ al. 2015, Thirty Meter Telescope Detailed Science Case: 2015
- Steidel etĀ al. (2011) Steidel, C.Ā C., BogosavljeviÄ, M., Shapley, A.Ā E., etĀ al. 2011, The Astrophysical Journal, 736, 160
- Steidel etĀ al. (2018) Steidel, C.Ā C., BogosavljeviÄ, M., Shapley, A.Ā E., etĀ al. 2018, The Astrophysical Journal, 869, 123
- Steidel etĀ al. (2010) Steidel, C.Ā C., Erb, D.Ā K., Shapley, A.Ā E., etĀ al. 2010, The Astrophysical Journal, 717, 289
- Steidel etĀ al. (2016) Steidel, C.Ā C., Strom, A.Ā L., Pettini, M., etĀ al. 2016, The Astrophysical Journal, 826, 159
- Sugahara etĀ al. (2017) Sugahara, Y., Ouchi, M., Lin, L., etĀ al. 2017, The Astrophysical Journal, 850, 51
- The Astropy Collaboration etĀ al. (2018) The Astropy Collaboration, Price-Whelan, A.Ā M., SipÅcz, B.Ā M., etĀ al. 2018, The Astronomical Journal, 156, 123
- The Astropy Collaboration etĀ al. (2013) The Astropy Collaboration, Robitaille, T.Ā P., Tollerud, E.Ā J., etĀ al. 2013, Astronomy & Astrophysics, 558, A33
- Tumlinson etĀ al. (2017) Tumlinson, J., Peeples, M.Ā S., & Werk, J.Ā K. 2017, Annual Review of Astronomy and Astrophysics, 55, 389, arXiv:1709.09180 [astro-ph]
- Turner etĀ al. (2014) Turner, M.Ā L., Schaye, J., Steidel, C.Ā C., Rudie, G.Ā C., & Strom, A.Ā L. 2014, Monthly Notices of the Royal Astronomical Society, 445, 794
- Tytler etĀ al. (2009) Tytler, D., Gleed, M., Melis, C., etĀ al. 2009
- UrbanoĀ Stawinski etĀ al. (2023) UrbanoĀ Stawinski, S.Ā M., Rubin, K. H.Ā R., Prochaska, J.Ā X., etĀ al. 2023, The Astrophysical Journal, 951, 135
- Urrutia etĀ al. (2019) Urrutia, T., Wisotzki, L., Kerutt, J., etĀ al. 2019, Astronomy & Astrophysics, 624, A141
- Vasan G.Ā C. etĀ al. (2024) Vasan G.Ā C., K., Jones, T., Shajib, A.Ā J., etĀ al. 2024, Spatially Resolved Galactic Winds at Cosmic Noon: Outflow Kinematics and Mass Loading in a Lensed Star-Forming Galaxy at $z=1.87$, arXiv:2402.00942 [astro-ph]
- Verhamme etĀ al. (2018) Verhamme, A., Garel, T., Ventou, E., etĀ al. 2018, Monthly Notices of the Royal Astronomical Society: Letters, 478, L60
- Virtanen etĀ al. (2020) Virtanen, P., Gommers, R., Oliphant, T.Ā E., etĀ al. 2020, Nature Methods, 17, 261
- Wang etĀ al. (2020) Wang, B., Heckman, T.Ā M., Zhu, G., & Norman, C.Ā A. 2020, The Astrophysical Journal, 894, 149
- Werk etĀ al. (2016) Werk, J.Ā K., Prochaska, J.Ā X., Cantalupo, S., etĀ al. 2016, The Astrophysical Journal, 833, 54
- Wisotzki etĀ al. (2016) Wisotzki, L., Bacon, R., Blaizot, J., etĀ al. 2016, Astronomy & Astrophysics, 587, A98
- Wisotzki etĀ al. (2018) Wisotzki, L., Bacon, R., Brinchmann, J., etĀ al. 2018, Nature, 562, 229
- Wolfe etĀ al. (1986) Wolfe, A.Ā M., Turnshek, D.Ā A., Smith, H.Ā E., & Cohen, R.Ā D. 1986, The Astrophysical Journal Supplement Series, 61, 249
- Xu etĀ al. (2023) Xu, Y., Ouchi, M., Nakajima, K., etĀ al. 2023, Stellar and AGN Feedback Probed with Outflows in JWST Galaxies at z=3-9: Implications of Frequent Nearly-Spherical Galactic Fountains, publication Title: arXiv e-prints ADS Bibcode: 2023arXiv231006614X
- Yuma etĀ al. (2017) Yuma, S., Ouchi, M., Drake, A.Ā B., etĀ al. 2017, The Astrophysical Journal, 841, 93
- Yuma etĀ al. (2013) Yuma, S., Ouchi, M., Drake, A.Ā B., etĀ al. 2013, The Astrophysical Journal, 779, 53
- Zabl et al. (2021) Zabl, J., Bouché, N. F., Wisotzki, L., et al. 2021, Monthly Notices of the Royal Astronomical Society, 507, 4294, arXiv:2105.14090 [astro-ph]